# DAS_CE_Supp

### v2

@@ Abstract
By deploying a large number of antennas with sub-half-wavelength spacing in a compact space, dense array systems (DASs) can fully unleash the multiplexing and diversity gains of limited apertures. To acquire these gains, accurate channel state information acquisition is necessary but challenging due to the large antenna numbers. To overcome this obstacle, this paper reveals that designing the observation matrix to exploit the high spatial correlation of DAS channels is crucial for realizing optimal/near-optimal channel estimation. Specifically, we prove that the observation matrix design for channel estimation is equivalent to a time-domain duality of point-to-point multiple-input multiple-output precoding, except for the change in the total power constraint on the precoding matrix to the pilot-wise discrete power constraint on the observation matrix. A novel ice-filling algorithm is proposed to design amplitude-and-phase controllable observation matrices, and a majorization-minimization algorithm is proposed to address the phase-only controllable case. Particularly, we prove that the ice-filling algorithm can be interpreted as a “quantized” water-filling algorithm, wherein the latter’s continuous power-allocation process is converted into the former’s discrete pilot-assignment process. To support the sub-optimality of the proposed designs, we provide comprehensive analyses on the achievable mean square errors and their asymptotic expressions. Finally, numerical results confirm that our proposed designs achieve the near-optimal channel estimation performance and outperform existing approaches significantly.
@@ I Introduction
Generally, DASs represent a series of array technologies that massive sub-wavelength antennas are densely arranged within a compact space. Unlike the conventional arrays with half-wavelength antenna spacing $\lambda/2$, the antenna spacing of DASs is much smaller, such as $\lambda/6$ [10], $\lambda/8$ [11], $\lambda/10$ [12], or even $\lambda/23$ [13]. With an increased number of sub-channels, DASs promise to achieve high array gains and fully exploit the multiplexing and diversity gains of limited apertures [9]. Besides, DASs can also reduce the grating lobes and provide high performance for large values of oblique angles of incidence [14]. Some works also report their advantages of super-bandwidth and super-directivity [15, 16, 17]. The typical DAS realizations include holographic multiple-input multiple-output (MIMO) [18], reconfigurable intelligent surfaces [19], fluid antenna systems [20, 21], graphene-based nano-antenna arrays [22, 23], radio-frequency lens [24], and so on. For example, in [10], massive sub-wavelength patches of $\lambda/6$-spacing are densely printed on a holographic MIMO surface to generate multiple beams flexibly. In [25], meta-elements sized of $\lambda/2.5\times\lambda/17.5$ are seamlessly integrated onto a feeding microstrip to achieve reconfigurable holographic surface (RHS) beam steering. In [13], an RIS composed of antennas sized of $\lambda/23\times\lambda/23$ is designed and fabricated for terahertz communications.
@@
The performance gains of antenna arrays are realized via the constructive beamformers enabled by their phase-shifters (PSs) and radio frequency (RF) chains. To realize effective beamforming and signal demodulation, the acquisition of accurate channel state information (CSI) is essential, particularly in situations where the number of RF chains is smaller than the number of massive antennas. Up to now, many estimators have been proposed to acquire the CSI of large antenna arrays. Their fundamental principle involves receiving pilot signals using observation matrices and recovering the channel by advanced estimation algorithms. For example, when the available pilot length is larger than the antenna number, the classical non-parametric algorithms are widely used, including least square (LS) estimator and minimum mean square error (MMSE) estimator. Leveraging the property of channel sparsity, compressed sensing (CS)-based channel estimators are widely studied to improve the estimation accuracy and reduce the pilot overhead [26, 27], such as the orthogonal matching pursuit (OMP)-based estimator [28, 29], the message passing (MP)-based estimator [30, 31, 32], and the gridless sparse signal reconstructor [33]. By training neural networks with a large amounts of channel data, the deep learning approaches are also utilized to realize data-driven channel estimators [34, 35, 36]. Besides, beam alignment techniques [37, 38, 39], including beam sweeping and hierarchical beam training, have also been widely explored to acquire the implicit CSI with low pilot overhead.
@@
Although most of existing channel estimators can be adopted in DASs, they fail to fully exploit the high spatial correlation of DAS channels, thereby leaving a remarkable performance gap from the optimal estimator. Specifically, this spatial correlation is attributed to the fact that, the extremely-dense deployment of DAS antennas significantly increases the similarity of radio waves impinging on antenna ports and aggravates the mutual-coupling effect between adjacent antenna circuits. From the mathematical perspective, this fact implies that the covariance matrices of DAS channels are no longer diagonal, but highly structured and under-determined. For diagonal covariance matrices, a uniform or random pilot-sensing strategy is sufficient to achieve the optimality of estimation [40, 41]. Thereby, in conventional channel estimations, the observation matrices for receiving the pilot signals are either generated randomly or set as predefined codebooks, such as the Discrete Fourier Transform (DFT) matrix and the identity matrix [26, 27, 28, 33, 30, 29, 31, 34, 42, 37, 38, 32, 36, 39]. In contrast, since the correlation matrix of DAS channels are full of structural features, it is believed that their observation matrices can be tightly aligned with these features in pilot transmission. This structured pilot-sensing process has a high potential to remarkably boost the channel estimation accuracy in DASs [43, 44, 45].
@@
To achieve this goal, our work represents the first attempt on designing and analyzing the optimal/near-optimal observation matrices in the DAS channel estimation. Our main contributions are summarized as follows.
@@
  • Ice-filling enabled observation matrix design: For practical DAS channel estimations, the continuous power-allocation by water-filling is not applicable as the pilot length is discrete and each pilot transmission has independent power constraint. To deal with this issue, we propose an ice-filling algorithm to design the amplitude-and-phase controllable observation matrices. This algorithm sequentially generates the optimal blocks of the observation matrix by maximizing the mutual information (MI) increment between two adjacent pilot transmissions. An important insight is that the ice-filling algorithm converts the power-allocation-process of the ideal water-filling algorithm into an eigenvector-assignment-process. We rigorously prove that, the continuous powers allocated by water-filling are quantized as the number of times that each eigenvector of the channel covariance is assigned for pilot transmission by ice-filling, with a quantization error smaller than one. The proven quantization nature thereby ensures the near-optimality of the proposed ice-filling algorithm.
@@ II System Model
Figure 1: An illustration of uplink channel estimation for an DAS.
@@
where $s_{q}$ is assumed to be 1 for all $q\in\{1,\cdots,Q\}$, ${\bf{W}}=\left[{\bf{w}}_{1},\cdots,{\bf{w}}_{Q}\right]$ stands for the observation matrix, and ${\bf z}:=\left[{\bf w}^{H}_{1}{\bf z}_{1},\cdots,{\bf w}^{H}_{Q}{\bf z}_{Q}\right]^{\rm T}$. As the power of ${\mathbf{w}}_{q}$ is normalized to 1, we have ${\bf z}\sim\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$. The objective of channel estimation is to recover the channel $\bf h$ from the received pilot $\bf y$. As a DAS possesses highly correlated channels across different antenna ports, we adopt Bayesian regression to recover ${\mathbf{h}}$. Suppose the channel is generated from a Gaussian process $\mathcal{CN}({\mathbf{0}}_{M},\bm{\Sigma}_{\bf h})$, where the covariance matrix $\bm{\Sigma}_{\bf h}\in\mathbb{C}^{M\times M}$, also called prior kernel, characterizes our prior knowledge to the channel correlation. The joint probability distribution of $[{\mathbf{h}}^{T},{\mathbf{y}}^{T}]^{T}$ satisfies
@@
Thereby, the estimated channel, $\hat{{\mathbf{h}}}$, is determined by the posterior mean of ${\mathbf{h}}$ given observation ${\mathbf{y}}$:
${{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}={{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}{{\bf{y}}}.$ (4)
The posterior covariance matrix of ${\mathbf{h}}$, also called posterior kernel, is accordingly derived as
@@
Notably, the posterier kernel, characterizing the channel estimation error, is a function of the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ and the observation matrix ${\mathbf{W}}$. Due to the extremely high spatial correlation exhibited by DAS channels, the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ would remarkably deviate from the identity matrix. This deviation indicates that aligning the observation matrix ${\mathbf{W}}$ with the channel’s subspace could be greatly helpful in reducing the estimation error. Motivated by this fact, this paper concentrates on the design of observation matrix ${\mathbf{W}}$, by considering both the amplitude-and-phase controllable and phase-only controllable analog combiners, to boost the channel estimation performance.
@@ III Mutual Information Maximization for Channel Estimation in DASs
In this section, we first presents a MIM based framework for designing observation matrices. Then, a water-filling-inspired observation-matrix-design is investigated as an ideal case of the proposed framework.
@@ III-A MIM based framework for observation matrix design
We adopt the MI between the received signal ${\mathbf{y}}$ and the channel ${\mathbf{h}}$ to evaluate the quality of the designed observation matrix. There are two reasons for using the MI as a metric. Firstly, MI tells the amount of information about the unknown channel ${\mathbf{h}}$ we can gain from the received signal ${\mathbf{y}}$. The second reason is that, according to the information-theoretic properties of Bayesian statistics from [46], the maximization of MI is asymptotically the minimization of Cramer Rao Bound (CRB).
@@
where $I(\cdot;\cdot)$ denotes the MI, $\det(\cdot)$ is the determinant of its argument, and $\mathcal{W}$ represents the feasible set of ${\mathbf{W}}$ for both amplitude-and-phase controllable and phase-only controllable analog combiners. It is notable that the observation matrix design problem in (6) resembles the well-know point-to-point MIMO precoding problem [47]. Comprehensive comparisons between these two methodologies are provided below.
  • Category of unknown data: Both MIMO precoding and observation matrix design in (6) aim to reduce the uncertainty between the received data, ${\mathbf{y}}$, and the unknown data. The former methodology regards the transmitted symbol vector, denoted by ${\mathbf{x}}$, as the unknown data. To facilitate MIMO precoding, multiple RF chains are employed to support an equal number of data streams. On the other hand, the unknown data to the latter methodology is the channel, ${\mathbf{h}}$. Our considered uplink channel estimation requires transmitting pilot signals over multiple timeslots to recover ${\mathbf{h}}$. In comparison, one can discover that the multiple RF chains in MIMO precoding are converted to the multi-timeslot pilot resources in channel estimation. Thereby, the design of the observation matrix in (6) can be viewed as a time-domain duality of point-to-point MIMO precoding.
@@ III-B Water-Filling Inspired Ideal Observation Matrix Design
Prior to addressing problem (6), we would like to consider an ideal (but practically unachievable) situation to establish an upper bound of the channel estimation performance. Specifically, we temporarily make the ideal assumption that the noise vector ${\mathbf{z}}$ in (2) is independent of the observation matrix ${\mathbf{W}}$ and its distribution is always $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$ regardless of the power of ${\mathbf{w}}_{q}$. Under this assumption, the pilot-wise power constraint, $\|{\mathbf{w}}_{q}\|^{2}_{2}=1$, can be relaxed to the total power constraint, $\|{\mathbf{W}}\|_{F}^{2}=\sum_{q=1}^{Q}\|{\mathbf{w}}_{q}\|^{2}_{2}=Q$. In this context, problem (6) is identical to the point-to-point MIMO precoding problem, which can be optimally solved using the water-filling principle [48]. It is crucial to emphasize again that this ideal situation is not practical because the noise vector ${\mathbf{z}}$ is inevitably amplified by observation matrix ${\mathbf{W}}$ during uplink channel estimation. This situation is only utilized to show the performance upper bound of channel estimation.
Let the eigenvalue decomposition of ${\bf\Sigma_{h}}$ be ${\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$, where $K$ is the rank of $\bf\Sigma_{h}$, ${\mathbf{U}}_{K}=[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$ encompasses the eigenvectors, and ${\bf\Lambda}_{K}={\rm diag}\{{\lambda}_{1},{\lambda}_{2},\cdots,{\lambda}_{K}\}$ represents the eigenvalues in a descending order. The water-filling algorithm dictates that the optimal observation matrix ${\mathbf{W}}$ is ${\mathbf{W}}={\mathbf{U}}_{K}{\mathbf{P}}$, where ${\mathbf{P}}\in\mathbb{R}^{K\times Q}$ represents the power allocation matrix. We assume that the pilot length $Q$ is larger than the rank $K$. Then power allocation matrix is expressed as ${\mathbf{P}}=[{\mathbf{P}}_{K},{\mathbf{0}}_{K\times(Q-K)}]$ with ${\mathbf{P}}_{K}={\rm diag}\{\sqrt{p_{1}},\sqrt{p_{2}},\cdots,\sqrt{p_{K}}\}$. The power $p_{k}$ allocated to each eigenvector is calculated by the water-filling principle:
@@
where the water-level $\beta$ is adjusted to meet the total power constraint: $\sum_{k=1}^{K}p_{k}=Q$.

### v3

@@ Abstract
By deploying a large number of antennas with sub-half-wavelength spacing in a compact space, dense array systems (DASs) can fully unleash the multiplexing and diversity gains of limited apertures. To acquire these gains, accurate channel state information acquisition is necessary but challenging due to the large antenna numbers. To overcome this obstacle, this paper reveals that designing the observation matrix to exploit the high spatial correlation of DAS channels is crucial for realizing near-optimal Bayesian channel estimation. Specifically, we prove that the observation matrix design for channel estimation is equivalent to a time-domain duality of point-to-point multiple-input multiple-output precoding, except for the change in the total power constraint on the precoding matrix to the pilot-wise discrete power constraint on the observation matrix. Inspired by Bayesian regression, a novel ice-filling algorithm is proposed to design amplitude-and-phase controllable observation matrices, and a majorization-minimization algorithm is proposed to address the phase-only controllable case. Particularly, we prove that the ice-filling algorithm can be interpreted as a “quantized” water-filling algorithm, wherein the latter’s continuous power-allocation process is converted into the former’s discrete pilot-assignment process. To support the near-optimality of the proposed designs, we provide comprehensive analyses on the achievable mean square errors and their asymptotic expressions. Finally, numerical results confirm that our proposed designs achieve the near-optimal channel estimation performance and outperform existing approaches significantly.
@@ I Introduction
Generally, DASs represent a series of array technologies that massive sub-wavelength antennas are densely arranged within a compact space. Unlike the conventional arrays with half-wavelength antenna spacing $\lambda/2$, the antenna spacing of DASs is much smaller, such as $\lambda/6$ [10], $\lambda/8$ [11], $\lambda/10$ [12], or even $\lambda/23$ [13]. With an increased number of sub-channels, DASs promise to achieve high array gains and fully exploit the multiplexing and diversity gains of limited apertures [9]. Besides, DASs can also reduce the grating lobes and provide high performance for large values of oblique angles of incidence [14]. The typical DAS realizations include holographic multiple-input multiple-output (MIMO) [15], reconfigurable intelligent surfaces [16], fluid antenna systems (FASs) [17, 18], graphene-based nano-antenna arrays [19, 20], radio-frequency lens [21], and so on. For example, in [10], massive sub-wavelength patches of $\lambda/6$-spacing are densely printed on a holographic MIMO surface to generate multiple beams flexibly. In [22], meta-elements sized of $\lambda/2.5\times\lambda/17.5$ are seamlessly integrated onto a feeding microstrip to achieve reconfigurable holographic surface (RHS) beam steering. In [13], a RIS composed of antennas sized of $\lambda/23\times\lambda/23$ is designed and fabricated for terahertz communications.
@@
The performance gains of antenna arrays are realized via the constructive beamformers enabled by their phase-shifters (PSs) and radio frequency (RF) chains. To realize effective beamforming, the acquisition of accurate channel state information (CSI) is essential, particularly in situations where the number of RF chains is smaller than the number of massive antennas. Up to now, many estimators have been proposed to acquire the CSI of large antenna arrays, and most of them can be adopted in DASs [23, 24, 25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37]. Their fundamental principle involves receiving pilot signals using observation matrices and recovering the channel by advanced estimation algorithms. For example, when the available pilot length is larger than the antenna number, the classical least square (LS) estimator is usually adopted. Leveraging the property of channel sparsity, compressed sensing (CS)-based channel estimators are widely studied to improve the estimation accuracy and reduce the pilot overhead [23, 24], such as the orthogonal matching pursuit (OMP)-based estimator [25, 26], the message passing (MP)-based estimator [27, 28, 29], and the gridless sparse signal reconstructor [30]. By training neural networks with a large amounts of channel data, the deep learning approaches are also utilized to realize both data-driven and model-driven channel estimators [31, 32, 33, 37]. Besides, beam alignment techniques [34, 35, 36], including beam sweeping and hierarchical beam training, have also been widely explored to acquire the implicit CSI with low pilot overhead.
@@
Although most of existing channel estimators can be adopted in DASs, they fail to fully exploit the high spatial correlation of DAS channels, thereby leaving a remarkable performance gap from the optimal estimator. Specifically, this spatial correlation is attributed to the fact that, the extremely-dense deployment of DAS antennas significantly increases the similarity of radio waves impinging on antenna ports and aggravates the mutual-coupling effect between adjacent antenna circuits. This fact makes the covariance matrices of DAS channels no longer diagonal but highly structured. For diagonal covariance matrices, the observation matrices for receiving pilot signals are either generated randomly or set as predefined codebooks, such as the Discrete Fourier Transform (DFT) matrix [23, 24, 25, 30, 32, 27, 26, 28, 31, 37, 34, 35, 29, 33, 36], which are sufficient to achieve the optimality of channel estimation [38, 39]. In contrast, since the correlation matrix of DAS channels are full of structural features, it is believed that their observation matrices can be tightly aligned with these features in pilot transmission. This structured pilot-sensing process has a high potential to remarkably boost the channel estimation accuracy in DASs [40, 41, 42]. For example, in [18], the structured channel covariance is utilized to determine the positions of fluid antennas during the pilot reception. Due to the strong channel correlation of FAS channels, the estimator in [18] can achieve higher accuracy than CS methods. Authors in [43] further extend the method in [18] to sparse channel scenarios. In [44], aiming at efficient CSI acquisition in holographic MIMO systems, the electromagnetic channel covariance is modeled and utilized to enable a Bayesian channel estimator. However, these existing estimators often rely on specific array structures or parametric channel models, which can neither be used to design the general observation matrix in DASs nor guarantee the performance to be optimal.
@@
To fill in this blank, our work represents the first attempt on designing and analyzing the near-optimal observation matrices in the DAS channel estimation. The classical Bayesian regression method is adopted as the channel estimator, i.e., the minimum-mean-square-error (MMSE) method, while our key contributions lie in the designed observation matrices to enhance Bayesian regression schemes, as summarized below.
@@
  • Ice-filling enabled observation matrix design: For practical DAS channel estimations, the continuous power-allocation by water-filling is not applicable as the pilot length is discrete and each pilot transmission has independent power constraint. To deal with this issue, we propose an ice-filling algorithm to design the amplitude-and-phase controllable observation matrices. Inspired by Bayesian regression, this algorithm sequentially generates the optimal blocks of the observation matrix by maximizing the mutual information (MI) increment between two adjacent pilot transmissions. An important insight is that the ice-filling algorithm converts the power-allocation-process of the ideal water-filling algorithm into an eigenvector-assignment-process. We rigorously prove that, the continuous powers allocated by water-filling are quantized as the number of times that each eigenvector of the channel covariance is assigned for pilot transmission by ice-filling, with a quantization error smaller than one. The proven quantization nature thereby ensures the near-optimality of the proposed ice-filling algorithm.
@@ II System Model
Figure 1: An illustration of uplink channel estimation for a DAS.
@@
where $s_{q}$ is assumed to be 1 for all $q\in\{1,\cdots,Q\}$, ${\bf{W}}=\left[{\bf{w}}_{1},\cdots,{\bf{w}}_{Q}\right]$ stands for the observation matrix, and ${\bf z}:=\left[{\bf w}^{H}_{1}{\bf z}_{1},\cdots,{\bf w}^{H}_{Q}{\bf z}_{Q}\right]^{\rm T}$. As the power of ${\mathbf{w}}_{q}$ is normalized to 1, we have ${\bf z}\sim\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$. The objective of channel estimation is to recover the channel $\bf h$ from the received pilot $\bf y$. As a DAS possesses highly correlated channels across different antenna ports, we adopt Bayesian regression to recover ${\mathbf{h}}$. Suppose the channel is generated from a Gaussian process $\mathcal{CN}({\mathbf{0}}_{M},\bm{\Sigma}_{\bf h})$, where the covariance matrix $\bm{\Sigma}_{\bf h}\in\mathbb{C}^{M\times M}$, also called prior kernel, characterizes our prior knowledge to the channel correlation. The kernel reflects the long-term statistical information of the covariance of channels, i.e., ${\bf\Sigma}={\mathsf{E}}({\bf h}{\bf h}^{H})$, which can keep unchanged for a long time. To obtain the prior kernel in practical MIMO systems, an ideal approach is to leverage the channel realizations generated by some standard models or real-world datasets. In addition, some online second-moment learning methods proposed in [45, 46] can also be used to obtain the prior kernel efficiently. Given the prior kernel, the joint probability distribution of $[{\mathbf{h}}^{T},{\mathbf{y}}^{T}]^{T}$ satisfies
@@
By invoking the property of the linear transform of Gaussian random vectors, the posterior mean and posterior covariance of ${\mathbf{h}}$ given observation ${\mathbf{y}}$ can be calculated as:
${{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}$ $={{\bf{\Sigma}}_{\bf{h}}}{{\bf{W}}}{\left({\bf W}^{H}{\bf{\Sigma}}_{\bf{h}}{\bf W}+\sigma^{2}{\bf I}_{Q}\right)^{-1}}{{\bf{y}}},$ (4)
@@
The optimal Bayesian estimation to the channel vector is exactly the posterior mean, i.e., $\hat{{\mathbf{h}}}={{\bm{\mu}}}_{{\mathbf{h}}|{\mathbf{y}}}$, which is also known as the MMSE estimator. Notably, the posterior covariance, ${\bf\Sigma_{{\mathbf{h}}|{\mathbf{y}}}}$, also called posterior kernel, characterizes the channel estimation error and is a function of the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ and the observation matrix ${\mathbf{W}}$. Due to the extremely high spatial correlation exhibited by DAS channels, the prior kernel ${{\bf{\Sigma}}_{\bf{h}}}$ would remarkably deviate from the identity matrix. This deviation indicates that aligning the observation matrix ${\mathbf{W}}$ with the kernel’s subspace could be greatly helpful in reducing the estimation error. Motivated by this fact, this paper concentrates on the design of observation matrix ${\mathbf{W}}$, by considering both the amplitude-and-phase controllable and phase-only controllable analog combiners, to boost the channel estimation performance.
@@ III Mutual Information Maximization for Channel Estimation in DASs
In this section, we first present a MIM based framework for designing observation matrices. Then, a water-filling-inspired observation-matrix-design is investigated as an ideal case of the proposed framework.
@@ III-A MIM based framework for observation matrix design
We adopt the MI between the received signal ${\mathbf{y}}$ and the channel ${\mathbf{h}}$ as the metric to evaluate the quality of the designed observation matrix. There are two reasons for using the MI as a metric. Firstly, MI tells the amount of information about the unknown channel ${\mathbf{h}}$ we can gain from the received signal ${\mathbf{y}}$. The second reason is that, according to the information-theoretic properties of Bayesian statistics from [47], the maximization of MI is asymptotically the minimization of Cramer Rao Bound (CRB).
@@
where $I(\cdot;\cdot)$ denotes the MI, $\det(\cdot)$ is the determinant of its argument, and $\mathcal{W}$ represents the feasible set of ${\mathbf{W}}$ for both amplitude-and-phase controllable and phase-only controllable analog combiners. It is notable that the observation matrix design problem in (6) resembles the well-know point-to-point MIMO precoding problem. Comprehensive comparisons between these two problems are provided below.
  • Category of unknown data: Both MIMO precoding and observation matrix design in (6) aim to reduce the uncertainty between the received data, ${\mathbf{y}}$, and the unknown data. The former regards the transmitted symbol vector, denoted by ${\mathbf{x}}$, as the unknown data. To facilitate MIMO precoding, multiple RF chains are employed to support an equal number of data streams. On the other hand, the unknown data to the latter is the channel, ${\mathbf{h}}$. Our considered uplink channel estimation requires transmitting pilot signals over multiple timeslots to recover ${\mathbf{h}}$. In comparison, one can discover that the multiple RF chains in MIMO precoding are converted to the multi-timeslot pilot resources in channel estimation. Thereby, the design of the observation matrix in (6) can be viewed as a time-domain duality of point-to-point MIMO precoding.
@@ III-B Water-filling Inspired Ideal Observation Matrix Design
Prior to addressing problem (6), we would like to consider an ideal (but practically unachievable) situation to establish an upper bound of (6). Specifically, we temporarily make the ideal assumption that the noise vector ${\mathbf{z}}$ in (2) is independent of the observation matrix ${\mathbf{W}}$ and its distribution is always $\mathcal{CN}({\mathbf{0}}_{Q},\sigma^{2}{\mathbf{I}}_{Q})$ regardless of the power of ${\mathbf{w}}_{q}$. In this context, the pilot-wise power constraint $\|{\mathbf{w}}_{q}\|^{2}_{2}=1$ can be relaxed to the total power constraint $\|{\mathbf{W}}\|_{F}^{2}=\sum_{q=1}^{Q}\|{\mathbf{w}}_{q}\|^{2}_{2}=Q$, and problem (6) becomes identical to the point-to-point MIMO precoding problem, which can be solved by water-filling [48].
Let the eigenvalue decomposition (EVD) of ${\bf\Sigma_{h}}$ be ${\mathbf{U}}_{K}{\bf\Lambda}_{K}{\mathbf{U}}_{K}^{H}$, where $K$ is the rank of $\bf\Sigma_{h}$, ${\mathbf{U}}_{K}=[{\mathbf{u}}_{1},{\mathbf{u}}_{2},\cdots,{\mathbf{u}}_{K}]$, and ${\bf\Lambda}_{K}={\rm diag}\{{\lambda}_{1},{\lambda}_{2},\cdots,{\lambda}_{K}\}$ with ${\lambda}_{1}\geq\cdots\geq\lambda_{K}>0$. The ideal observation matrix ${\mathbf{W}}^{\rm Ideal}$ is given as ${\mathbf{W}}^{\rm Ideal}={\mathbf{U}}_{K}{\mathbf{P}}$. The power allocation matrix ${\mathbf{P}}\in\mathbb{R}^{K\times Q}$ is expressed as ${\mathbf{P}}=[{\mathbf{P}}_{K},{\mathbf{0}}_{K\times(Q-K)}]$ with ${\mathbf{P}}_{K}={\rm diag}\{\sqrt{p_{1}},\cdots,\sqrt{p_{K}}\}$. The power $p_{k}$ allocated to each eigenvector is calculated by the water-filling principle:
@@
where $(x)^{+}\overset{\Delta}{=}\max\{x,0\}$, and the water-level $\beta$ is properly selected to meet the power constraint: $\sum_{k=1}^{K}p_{k}=Q$.